OpenStudy (anonymous):

(ALGEBRA!) Joe's average score on 3 tests was 79. His average for his next 4 tests were 78.5. There are 2 tests left in the semester and Joe needs to raise his average to at least 80. What must he average on the next 2 tests?

OpenStudy (anonymous):

HELP PLEASE!

OpenStudy (anonymous):

@nabizag help?

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

let x be the total for the first three y = total of the next 4 tests z = total of the next 2 tests.. okay so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@nincompoop help?

OpenStudy (lgbasallote):

sorry i got lagged up. anyway the final average would be (x+y+z)/9 >= 80 correct?

OpenStudy (anonymous):

80%?

OpenStudy (lgbasallote):

uhh...no...im just interpreting your problem into algebra...it's not the answer yet

OpenStudy (lgbasallote):

your question said the final average is 80% so that's what i wrote

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

anyway (x+ y + z)/9 \(\ge\) 80 now cross multiply..what do you get?

OpenStudy (anonymous):

do i add 78.5, 80, and 79?

OpenStudy (anonymous):

then divide by 9?

OpenStudy (lgbasallote):

no..

OpenStudy (lgbasallote):

78.5 and 79 are averages...x y and z are totals

OpenStudy (lgbasallote):

do you see what im planning to do? basically what you're aiming to do is: score for first quiz + score of second quiz + score of third quiz + ....+ score of 9th quiz then divide it all by 9 and the result should be greater than or equal to 80 make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

92.61 I THINK... I shall now try and figure out how the hell I came to this figure......

OpenStudy (lgbasallote):

good. so you have \[\frac {x + y + z}9 \ge 80\] cross multiply..what do you get?

OpenStudy (anonymous):

26.3 less than or equal to 80 then subtract the 26.3?

OpenStudy (anonymous):

Actually no, Im pretty sure im wrong...

OpenStudy (lgbasallote):

why do you have 26.3?

OpenStudy (anonymous):

I divided the answer i got of all 3 numbers by 9

OpenStudy (lgbasallote):

just focus on the equation im just showing you right now \[\frac{x+y+z}9 \ge 80\] cross multiply don't put in any additional information

OpenStudy (anonymous):

okay I got 237.7 greater than or equal to 80

OpenStudy (anonymous):

How do i cross multiple that, Im not sure how too.

OpenStudy (lgbasallote):

why are you getting numbers? i wrote letters and they come out numbers?

OpenStudy (lgbasallote):

you just multiply both sides by 9. nothing more

OpenStudy (anonymous):

okay so i got 9x+9y+18>=720

OpenStudy (lgbasallote):

if you multiply (x+y+x)/9 by 9..the 9 just cancels out..so it should be \[x + y + z \ge 720\] got it?

OpenStudy (anonymous):

okay yes

OpenStudy (lgbasallote):

good so remember that equation..we'll come back to that later. for now..we gve x and y some values

OpenStudy (lgbasallote):

it says the average of the first 3 tests is 79..so you can equate this as \[\frac x3 = 79\] right? since x is the total of the first 3 tests

OpenStudy (anonymous):

The answer is definitely 84.5

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so now solve for x by multiplying both sides by 3

OpenStudy (anonymous):

x=237

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

now it says the average of the next 4 quizzes is 78.5 so \[\frac y4 = 78.5\] solve for y

OpenStudy (anonymous):

y=314

OpenStudy (lgbasallote):

right. so you have x = 237 and y = 314 now substitute that into x +y + z > = 720 then solve for z

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

what do you get?

OpenStudy (anonymous):

z>169

OpenStudy (lgbasallote):

close.. z \(\ge\) 169 since z is the total of the next two tests, to get the average, divide 169 by 2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

84.5

OpenStudy (lgbasallote):

right. so he needs at least 84.5 average in the next two tests

OpenStudy (anonymous):

thank you!

OpenStudy (lgbasallote):

indeed it was

OpenStudy (anonymous):

I concur! ;)